OpenStudy (anonymous):

MEDAL!!!:) PLEASE SOMEBODY HELP ME! In 1985, there were 285 cell phone subscribers in Cleveland, Ohio. In 1988, there were 1,527 and in 1994, there were 43,871. Assuming that this is exponential growth, find the exponential regression equation and use it to find the number of cell phone subscribers in 1990. Round to the nearer hundred. A. 4,700 B. 6,600 C. 6,800 D. 7,000

OpenStudy (agent0smith):

You could do this the way we did the other exponential one, though it's clear now you weren't supposed to do it that way (exponential regression equation) you'd have to use a calculator. Since you said you don't have one, follow the method we used last time.

OpenStudy (anonymous):

Okay I used what I did last time and got and answer of 6,600 is that correct, or no?

OpenStudy (agent0smith):

1985, there were 285 - let's use that as Ao \[\Large A = 285 e^{rt}\] In 1988, there were 1,527... 3 years after '85, so t=3 \[\Large 1527 = 285 e^{3r}\]find r. then, plug in r, and 1990 is 5 years after that, so t= 5\[\Large A = 285 e^{5r}\] and find A

OpenStudy (agent0smith):

I didn't get 6600

OpenStudy (anonymous):

hmmm how would I find A? divide?

OpenStudy (agent0smith):

Once you have r, plug it in and calculate. Find r first.

OpenStudy (anonymous):

ok hold on quick

OpenStudy (anonymous):

okay I get 1=5 for r

OpenStudy (anonymous):

1/5

OpenStudy (anonymous):

so now what do i do

OpenStudy (agent0smith):

r=1/5 isn't correct. Do you know how to take logs of both sides? \[\Large 1527 = 285 e^{3r}\] divide both sides by 285:\[\Large 5.3578 = e^{3r}\]now take logs of both sides

OpenStudy (anonymous):

wait so this is what I do? 1527=285e^3r/285 and 5.3578=e^3r/285

OpenStudy (agent0smith):

Yes, then you need to take logs of both sides

OpenStudy (agent0smith):

Except 5.3578=e^3r no 285 now

OpenStudy (anonymous):

okay, it seems as if my answer is leading to 4,700

OpenStudy (anonymous):

that would be correct right?

OpenStudy (anonymous):

hey quick yes or no question: What is the value of c in the quadratic regression equation for these points: (-3, 1.2), (1, 0.4), (5, 1.6), (11, 2.6)? A. -0.01 B. 0.16 C. 0.83 D. 0.86 wouldnt my answer be D

OpenStudy (agent0smith):

That one idk... i think your last question was a quadratic though, i'd do whatever you did in the last question.

OpenStudy (anonymous):

okay but 4,700 is the right answer.

OpenStudy (agent0smith):

Looks that way, based on my work too... but again, we didn't use exponential regression